### v1

# Physics-Informed Detection of Friction Anomalies in Satellite Reaction Wheels

Abstract
As the number of satellites in orbit has increased exponentially in recent years, ensuring their correct functionality has started to require automated methods to decrease human workload. In this work, we present an algorithm that analyzes the on-board data related to friction from the Reaction Wheel Assemblies (RWA) of a satellite and determines their operating status, distinguishing between nominal status and several possible anomalies that require preventive measures to be taken. The algorithm first uses a model based on hybrid systems theory to extract the information relevant to the problem. The extraction process combines techniques in changepoint detection, dynamic programming, and maximum likelihood in a structured way. A classifier then uses the extracted information to determine the status of the RWA. This last classifier has been previously trained with a labelled dataset produced by a high-fidelity simulator, comprised for the most part of nominal data. The final algorithm combines model-based and data-based approaches to obtain satisfactory results with an accuracy around $\mathbf{95\boldsymbol{\%}}$.
Index Terms:
satellite, reaction wheel assembly, friction, anomaly detection, physics-informed
I Introduction
As satellites orbiting Earth have become more frequent during the last decades [1], the need to maintain their correct functioning has become a major concern of space agencies around the world. Being in space during their whole lifetime, satellite repairs are, in general, unfeasible, so failures usually lead to the termination of the mission. Even when a repair is possible, it tends to be very expensive, as was the case of the installation of a new camera for the Hubble telescope lens, which required an entire mission [2]. This makes the capability of detecting impending failures a fundamental component of their monitoring.
Among the components that can fail in a satellite, the Reaction Wheel Assemblies (RWAs) are of particular interest. RWAs are a fundamental component of the Attitude and Orbit Control System (AOCS) of the satellite. Their purpose is to exert a torque on the body of the satellite to control its attitude. Rather than using action-reaction principles as a Reaction Control System (RCS) does, an RWA uses angular momentum conservation to achieve greater accuracy with no propellant consumption, which is highly advantageous for star tracking and long-duration missions.
The operating principle of RWAs is fairly simple. Inside the assembly, a flywheel is kept rotating permanently. When there is a need to rotate the satellite in one direction, the disc is accelerated in the opposite direction using an electric motor. Due to the conservation of angular momentum in space, this produces the desired torque in the main body of the satellite [3].
Figure 1: Cross-section of a Reaction Wheel Assembly, adapted from [4].
To get complete control of the three axes of rotation of the satellite, at least 3 RWAs need to be mounted in a satellite. However, to achieve redundancy and some flexibility in the speed of rotation of the discs, it is common to install 4 RWAs instead [5].
The RWA is, then, a mechanical component in almost permanent motion, making it more prone to failure than other fixed satellite parts. Mechanical degradation, vibrations, and wear can occur in this element, compromising its integrity [6]. Particularly, the bearings, where contact between the rotating disc and the non-moving satellite body takes place, are home to several complex phenomena involving elastohydrodynamics (interaction between a moving fluid, such as lubricant, and deformable solid material). These phenomena cause sudden changes in friction at the bearings that make the wheel modeling particularly challenging. A failure of a RWA might imply losing all device functionality and can seriously compromise the mission’s success. Such has been the case in space missions Kepler [7], Dawn [8], and SDO [9], in which the capabilities of the satellites were compromised after RWA failures due to the loss of authority over the torque of the satellite [10, 9].
Although there have been attempts at explaining the causes of such failures [11], the complexity in the dynamics of the bearings makes it hard to determine the condition of the component, and the limited number of sensors in the RWA implies that a complete monitoring of the situation is impossible. Even though the nature of these failures remains unknown, recent studies have shown that they are usually preceded by changes in the behavior of the friction at the bearings, which can be observed using on-board sensors commonly installed in RWAs [12]. Such changes of behavior (that is, anomalies) are usually caused by changes in the properties of the RWA, such as alterations of the lubricant or an increase in the contact surface in the bearings.
It is then of interest to detect these changes in the behavior, as they can be used to anticipate possible future failures of the wheel and/or take mitigative measures to improve the health of the RWA [4]. However, not all changes imply a future failure, so it is also necessary to distinguish between different types of anomalous behavior to provide the appropriate warnings. Therefore, for this work, it is not enough to perform anomaly detection, but categorization should also be considered.
In this paper, we develop a method that analyzes the on-board data collected from an RWA during a window of time of a few hours and determines whether an anomaly is present in the RWA and, if positive, which one among those known, with the entire process carried out in the satellite thanks to its low computation cost. We propose the following contributions:
  • We derive a novel model for friction systems that incorporates sudden variations of friction due to quick changes in the internal geometry of the system. The model can be implemented without knowledge of the actual causes of these phenomena, but rather just their statistical properties, which can be obtained experimentally.
  • We provide an end-to-end algorithm that takes a window of a few hours of measurements by on-board sensors to determine whether the observed friction characteristics of the RWA belong to nominal or anomalous conditions. The method combines the use of the model together with data-driven approaches to provide an accurate diagnosis.
  • We include in the approach a modification to the classification of changepoints that accounts for the possibility of false negatives. This makes the algorithm robust to incorrectly detected changepoints, an unavoidable phenomenon due to the stochastic nature of the signal measured.
  • We thoroughly validate the complete algorithm and each of its components using high-fidelity simulation data from Thales Alenia Space (TAS), with no knowledge of the model used to generate it.
Literature review
Anomaly and/or fault detection in satellites has received considerable interest in recent years, primarily due to their proliferation and the need for constant monitoring. Research done in this direction has attempted to produce algorithms capable of analyzing all the information coming from the satellite to determine a wide variety of possibly faulty conditions. The high availability of data and the lack of models for the complete satellite has led to almost exclusively data-driven approaches, using tools such as Support Vector Machines (SVM) [13], Relevance Vector Machines (RVM) [14], Principal Component Analysis (PCA) [15], or LSTM networks [16]. These ideas have led to algorithms used in actual space missions, like Orca [17] in the International Space Station and the IMS tool [18] in the Space shuttle.
Satellite-specific methods are very versatile as they can catch a wide range of possible malfunctions in the vehicle. However, they consider the vehicle as a whole, ignoring the intricacies of each component. This makes them less sensitive to subtle changes in the behavior of any of these individual components.
It is then reasonable to find detection methods specific to the component of interest. Focusing on RWA, the simplicity and generally well-understood dynamics of these devices have allowed the development of more specialized techniques capable of leveraging expert knowledge. This has led to extremely detailed models to predict vibrations in the RWA [19, 20]. Although accurate, such models are unsuitable for anomaly detection during space missions, as they usually involve signals not available outside ground tests.
The problem of fault and anomaly detection for RWAs has been analyzed mainly through machine learning [21, 22]. Although these methods have been effective in their purpose, they consider the behavior of the electric motor rather than focusing on the friction torque, so they are not so precise in determining abnormal behavior of the latter. This was done only recently using unsupervised machine learning [23], but considering only normal and abnormal behavior, that is, without isolation of the anomaly.
Closely related to RWAs, rotary machinery analysis is a richer field of literature. Although purely model-based approaches have been proposed in the past [24], they are usually focused on alterations in the vibration patterns of the machine. Most of the research is based on data-driven techniques, capable of assessing a wider variety of faulty conditions usually related to problems in the bearings, as highlighted in the survey [25].
Most of the rotary machinery analysis techniques follow a common approach: first, some signals are measured from the machine (usually displacement or acceleration), then features are extracted (statistical ones like mean and variance, or frequency-related like Power Spectral Density), and finally, some machine learning methods are used to determine the status of the machine from these features. Such methods include nearest neighbor [26], neural networks [27], SVM [28], linear regressions [29], or Broad Learning Systems (BLS) [30], among others.
These methods have achieved important improvements in recent years by incorporating knowledge of a model of the machine, leading to the family of methods known as hybrid methods [31]. A popular technique is to fuse these two sources of knowledge through Physics Informed Machine Learning (PIML), which has experienced a rapid development in the last years [32].
Although evident progress has been achieved as new features and classification methods have been developed, their focus on the vibrational behavior of the machine makes them unsuitable for satellite applications, where this behavior is not of much interest. Regarding friction-related research, most of the interest is on the low-speed behavior, where complex phenomena like the Stribeck curve or stick-slip occur [33, 34, 35]. Research on switching behavior in friction has also experienced some attention, but it is restricted to deterministic systems[36].
In conclusion, there has not been any work in the study of friction systems (particularly RWAs) with random jumping friction behavior, which poses unique challenges for detecting anomalies. This article aims to fill the existing gap in the literature by providing a hybrid model-data-based anomaly detection and isolation algorithm.
II Problem Formulation
We study the problem of determining if the friction torque of an RWA has an anomalous behavior based on a window of measurements taken by its sensor system at fixed time intervals. Each measurement at time $t_{k}$ (where $t\in\mathbb{R}$, and $k\in\mathbb{N}$ is the index of the measurement) consists of the spin rate $\omega_{k}$ of the RWA, the voltage $V_{k}$, and the current $I_{k}$ in the motor that accelerates (or deccelerates) the RWA. Using conservation of angular momentum, energy equations at the motor and filtering techniques, we can obtain an unbiased estimation of the friction torque $f_{k}$ at the bearings of the RWA using an estimator:
$J\dot{\omega}$ $=T_{\text{motor}}+f=K_{T}I+f$ (1a)
$\hat{f}_{k}$ $=J\frac{\omega_{k+1}-\omega_{k-1}}{t_{k+1}-t_{k-1}}-K_{T}I_{k}$ (1b)
where $T_{\text{motor}}$ is the torque produced by the motor, $J$ is the wheel inertia and $K_{T}$ is the torque constant of the motor. Additionally, the estimated $\hat{f}_{k}$ can be filtered to remove the effect of noise. In this paper, we do not consider the details of this estimation process, but rather assume that $\hat{f}_{k}=f_{k}+v_{k}$, where $v_{k}$ is random noise with zero mean, assuming the estimation is unbiased.
Since we are interested only in the friction characteristics of the RWA, the input to our algorithm is a sequence of pairs of spin rate and estimated friction torque obtained in a window of time:
$X=\{(\omega_{k},\hat{f}_{k})\}_{k=1}^{N}.$ (2)
Both the spin rate and the estimated friction torque are normalized using average values in historical data, so the variables take values of order $1$. To ease the reading of the paper, a table with all the indices used in the paper is given in Table I. Subscripts and superscripts that are part of the name of the variable are written upright (like $N_{\mathrm{cp}}$) and those that are indices are written in italics (like $\omega_{k}$).
Symbol Meaning Range
$\ell$ Datapoint in dataset [1, $N_{\mathrm{D}}$]
$k$ Measurement in datapoint $1,N_{\mathrm{m}}]$
$i$ Changepoint $1,N_{\mathrm{cp}}]$
$s$ Switching system $1,N_{\mathrm{s}}]$
TABLE I: Indexes used in the paper
In this paper, we want to detect anomalies, defined as abnormal behavior in the friction torque of the RWA. Several possible anomalies are considered, each due to different phenomena occurring in the RWA. For a window of measurements $X$, any anomaly can be active or inactive: the status of the RWA is defined as a vector of boolean values specifying whether each anomaly is active or inactive: $\Theta=\{0,1\}^{N_{\mathrm{anom}}}$, where $N_{\mathrm{c}}$ is the number of anomalies considered.
Anomalies do not have an explicit mathematical description. Instead, our knowledge about them comes from a labeled dataset $D$ with windows of measurements annotated with the anomalies affecting the RWA that produced such measurements:
$D=\{(X_{\ell},\theta_{\ell})\}_{\ell=1}^{N_{\mathrm{D}}},$ (3)
where $\theta_{\ell}\in\Theta$. This dataset can be obtained from ground tests, in-orbit experience, or hight-fidelity simulators, for example.
To illustrate the shape of the datapoints in such a dataset, and how they might differ depending on the anomaly, we show three examples taken from the TAS dataset in Figure 2. Some early observations indicate that, even in the nominal case, there are sudden variations of the friction independent of the spin rate (for example, the one close to measurement $4000$ in the nominal case). Sudden changes in the bearings cause such variations due to, for example, oil accumulations in small pockets. Although their presence does not necessarily imply an anomaly in the RWA, they must be considered, since excessive friction changes due to these variations are usually concerning. A model that captures the behavior of an RWA must, therefore, model these sudden friction changes.
Regarding anomalies, Anomaly 1 produces a constant increase of friction torque at all spin rates, while Anomaly 2 increases friction torque proportionally with the spin rate. Additional anomalies relate to friction differences produced in the sudden events described in the previous paragraph.
Figure 2: Examples of runs labeled with different anomalies. For each case, the anomalous run is shown in red while a nominal run is plotted in blue to help compare them.
Based on these observations and the given dataset from TAS, we state the following problem.
Problem Statement: Create a model that captures the behavior observed in the TAS dataset, and use it to produce an algorithm that can determine the anomaly affecting an RWA based on a window of spin rate and friction torque measurements.
III RWA Modeling
To solve the problem formulated in the previous section, we propose a hybrid approach to overcome the usual limitation of purely data-based and model-based methods. In this section, we derive the model used for the model-based part, developed from a combination of dataset observation, knowledge of the RWA mechanism, and literature on the topic.
III-A Friction Model
Since the data available for the detection of anomalies consists of spin rate and friction torque signals, we discard the dynamical part shown in (1) and model the RWA as a friction system, that is, an algebraic relation between friction and spin rate:
$f_{k}$ $=f_{k}^{\mathrm{d}}\text{sign}(\omega_{k})+f^{\mathrm{v}}\omega_{k}+v_{k}$ (4a)
$f_{k}^{\mathrm{d}}$ $=f^{\bar{\mathrm{d}}}+\sum_{s=1}^{N_{\textrm{s}}}f_{k}^{s}$ (4b)
where (4a) corresponds to the classical dry and viscous friction with additive Gaussian noise $v$. The dry and viscous friction coefficients $f_{k}^{\mathrm{d}}$ and $f^{\mathrm{v}}$, respectively, are at the basis of any friction model [37], but note here that the dry coefficient is not constant and can change with time.
Each component of (4a) models a different friction factor. Starting from the end, the Gaussian noise $v$ has three sources: randomness in the motion of the bearing and the contact inside it, sensor noise, and inaccuracies in the friction torque estimation. Among all possible noise models available, the Gaussian one was chosen for its theoretical properties and its popularity in the literature [38, 39, 40].
The viscous friction, on the other hand, is produced by the lubricant inside the bearing [41]. This parameter is known to be affected by the temperature and the pressure of the lubricant, as well as its age [42]. However, these variations take place in longer time intervals than we analyze; therefore, we can assume this coefficient to be constant during the periods of time analyzed with the algorithm. Note that, although constant, this coefficient can still be affected by an anomaly that increases its value.
Finally, the dry friction models the contact between solid surfaces. In the bearings of an RWA, this contact occurs between the the inner and outer races of the bearing and the balls, and between the balls themselves [20]. The contact surface at every instant is not necessarily constant, which can lead to changes in the dry friction coefficient. To represent these changes, we first define in (4b) the dry friction coefficient as the sum of several terms, the first of which is the constant baseline dry friction $f^{\bar{d}}$ that remains constant at all times.
The remaining components $f_{k}^{s}$ are variable in time, and capture the effect of variations in the contact surface of the bearings. The evolution in time of each of these components is described by what we refer to as a Friction Switching System (FSS). The motivation behind using several FSS in the model of a single RWA is to capture independent phenomena with different properties. Such phenomena can range from concentrations of pockets of oil (well reported in the literature [41] to displacements of the different components of the bearing (observed in the past but not well characterized so far [6]), which may have very different time scales. By modeling them with different FSS, we can also distinguish anomalies affecting one or another phenomenon, enhancing the accuracy of the categorization of anomalies. An example of using several FSS in a single RWA model is given in Example 1.
III-B Friction Switching System
We define a Friction Switching System (FSS) as a hidden Markov Model (HMM) [43] with state $(q,f,\tau)$ where $q\in\llbracket 1,q_{\text{max}}\rrbracket$, $f\in\mathbb{R}$ and $\tau\in\mathbb{N}^{\geq 0}$, where $\mathbb{R}$ is the set of real numbers and $\mathbb{N}^{\geq 0}$ is the set of natural numbers including 0.
An FSS represents a particular phenomenon inside the bearing of the RWA, such as the distribution of lubricant, the relative position of the inner and outer races, or any other aspect that may affect the dry friction. In these systems, the variable $q$ represents the current configuration of the system, $\tau$ is how long the system has been in that configuration, and $f$ is the friction generated by the system. Note that the configuration only takes discrete values.
These systems are indexed with the variable $s$, and the evolution of a system $s$ is relatively simple. At every step, the system may jump to a contiguous configuration $q$ with $P_{s}^{\mathrm{t}}(q,\tau)$ or stay the same. If it does not jump, the configuration does not change, and $\tau$, the number of steps since the last jump, increases by one. If it jumps, the configuration $q$ and the friction $f$ produced by the FSS change, and $\tau$ goes back to 0. A visualization of a system FSS can be seen in Figure 3, and the dynamics are given formally by
$\left[\begin{matrix}q_{k+1}\\ f_{k+1}\\ \tau_{k+1}\end{matrix}\right]$ $=\begin{cases}\left[\begin{matrix}q_{k}&f_{k}&\tau_{k}+1\end{matrix}\right]^{\top}&\text{with prob. }P_{s}^{\mathrm{t}}(q_{k},\tau_{k}),\\ \left[\begin{matrix}q^{\prime}&f^{\prime}&0\end{matrix}\right]^{\top}&\text{otherwise}.\end{cases}$ (5)
Two probability functions affect the behavior of the FSS. The first is the transition probability $P_{s}^{t}(q,\tau)$, which depends both on the current configuration and on how long the system has spent in it. By including $q$, we can consider systems that spend different lengths of time in various configurations, and by including $\tau$, we are not limited to exponential residence times, as it is common in stochastic Markov chains [43].
The second probability function is $P_{s}^{\mathrm{qf}}(q^{\prime},f^{\prime})$, which determines the new configuration and friction ($q^{\prime}$ and $f^{\prime}$ in (5)) the system takes after it has transitioned. Since the friction depends only on the new configuration, this can be written as
$(q^{\prime},f^{\prime})\sim P_{s}^{\mathrm{qf}}(q^{\prime},f^{\prime}|q)=P_{s}^{\mathrm{f}}(f^{\prime}|q^{\prime})P_{s}^{\mathrm{q}}(q^{\prime}|q).$ (6)
We set some assumptions on $P_{s}^{\mathrm{f}}(f^{\prime}|q^{\prime})$ and $P_{s}^{\mathrm{q}}(q^{\prime}|q)$. For the latter, the FSS can only jump between contiguous states, that is, $P_{s}^{\mathrm{q}}(q^{\prime}|q)=0$ if $|q^{\prime}-q|>1$. For the friction: the probability functions associated with different $q$ can not overlap, and the range of values of $f^{\prime}$ must increase with increasing $q$. Figure 3 shows an example of valid friction probabilities on the right side.
To characterize a FSS, we need four elements: its number of configurations $q_{\text{max}}$, its transition probability function $P_{s}^{\mathrm{q}}(q^{\prime}|q)$, its jumping probability function $P_{s}^{\mathrm{t}}(q,\tau)$ for every configuration $q$ and the friction probability function $P_{s}^{\mathrm{f}}(f|q)$, also for every $q$. An example of the friction produced by an RWA modeled with two FSS is shown in Figure 4, where its similarity to the examples shown in Figure 2 can be noted.
Figure 3: Evolution of the state and friction of a switching system. The left nodes show the different configurations an FSS can take (3 in this example) and the possible transitions between them. The right side shows the friction distributions of each configuration, together with a possible trajectory of the friction value generated by the FSS as it jumps between configurations.
Example 1 (Short term friction changes)
We want to model sudden and short-lived increases of friction. This system has two configurations: inactive ($q=0$, no friction increase) and active ($q=1$, increased friction). The system stays inactive most of the time, getting activated only for a few steps before jumping back. Such behavior have been hypothesized to be caused by accumulations of oil pockets in the RWA bearings [41], and can be modeled with the functions
$P_{1}^{\mathrm{f}}(f^{\prime}|q^{\prime})$ $=\begin{cases}10\quad&\text{if }0.3q^{\prime}<f<0.3q^{\prime}+0.1\\ 0\quad&\text{otherwise}\end{cases}$ (7a)
$P_{1}^{\mathrm{t}}(\tau|q)$ $=\begin{cases}1/(200-\tau)\quad&\text{if }q=0\\ 1/(2000-\tau)\quad&\text{if }q=1\end{cases}$ (7b)
Example 2 (Longer term friction changes)
We want to model long-term variations of the friction switching between three different configurations, with the same transition probabilities. This can be due to, for example, changes in the angle of the cage inside the bearing. The probability functions are given by:
$P_{2}^{\mathrm{f}}(f^{\prime}|q^{\prime})$ $=\begin{cases}1.5\quad&\text{if }0.5q^{\prime}<f^{\prime}<0.5q^{\prime}+0.4\\ 0\quad&\text{otherwise}\end{cases}$ (8a)
$P_{2}^{\mathrm{t}}(x,\tau)$ $=\begin{cases}\frac{1}{30000-\tau}\quad&\text{if }x>10000,\\ 0\quad&\text{otherwise}.\end{cases}$ (8b)
We show a simulation of the friction system proposed in Equations (4a) and (4b) with the switching system using the probability functions described above and the parameters $f^{\bar{\mathrm{d}}}=1$ and $f^{\mathrm{v}}=1$. The evolution of the friction components in time, along with the total dry friction coefficient, is shown in Figure 4.
Figure 4: Visualization of the two friction systems described in Examples 1 and 2. The first three plots show the evolution of the three state variables of the system. The fourth plot shows the resulting total dry friction, and the fifth the total friction after adding a viscous component.
Remark 1
In the paper, we assume knowledge of $q_{\text{max}}$, $P_{s}^{\mathrm{t}}(x,\tau)$ and $P_{s}^{\mathrm{q}}(q^{\prime}|q)$ for every FSS in the RWA we are analyzing. Such knowledge may come from historical data, expert knowledge, or both. On the other hand, we do not know the friction probability function $P_{s}^{\mathrm{f}}(f^{\prime}|q^{\prime})$, which might be affected by an anomaly.
III-C Anomaly Modeling
To conclude this section, we define how the possible anomalies in the RWA are represented in the proposed model. As mentioned in Section I, there are several possible anomalies, and all of them affect the resulting friction experienced in the bearing of the RWA.
We consider each anomaly to relate to one and only one component of friction. For instance, the dry friction anomaly $\theta^{\mathrm{d}}$ affects the base dry friction coefficient, while the viscous friction anomaly $\theta^{\mathrm{v}}$ affects the viscous friction coefficient.
For the case of the switching systems, each FSS present in the RWA can be affected by one anomaly, represented by $\theta_{s}$, where $s$ is the index of the affected FSS. In this model, we consider that the effect of the anomaly manifests itself only in the friction distribution function, that is, in the values of the friction the system produces for every configuration. This relationship can be expressed mathematically as:
$f^{\bar{\mathrm{d}}}=f^{\bar{\mathrm{d}}}(\theta^{\mathrm{d}}),\quad f^{\mathrm{v}}=f^{\mathrm{v}}(\theta^{\mathrm{v}}),$ (9a)
$P_{s}^{\mathrm{f}}(f^{*}|x^{*})=P_{s}^{\mathrm{f}}(f^{*}|x^{*};\theta^{s}).$ (9b)
Although the probability functions are unknown, they can be estimated using a dataset $\mathcal{D}$, hence the need for a hybrid approach.
IV Algorithm Description
The proposed anomaly detection algorithm uses the model developed in Section III to analyze the window of measurements under investigation and correctly diagnose the state of the RWA. As explained in Section III, the possible anomalies of the RWA affect the values of friction that each component (base dry, viscous and the FSSs) produces. The objective, then, is to extract the relevant information from the data to do the classification.
The main challenge in the model presented in the previous section is the presence of several FSSs, each producing sudden changes in friction at random moments. It is then fundamental to keep track of these changes to detect and categorize the different anomalies.
For this purpose, the algorithm comprises four stages, each with a clear responsibility in the complete process. The first stage, changepoint detection, determines the time-steps in which an abrupt friction change has occurred. The second stage is the friction estimation, which determines the dry and viscous friction coefficients before and after every changepoint, that is, $f_{k}^{\mathrm{d}}$ and $f^{\mathrm{v}}$ in (4a). The different friction coefficients can then be used to calculate how much the friction changes at every changepoint. The third stage is the maximum likelihood assignment, which determines which FSS has produced each changepoint by maximizing the likelihood of the observed data. With this information we can distinguish the individual contributions $f^{\bar{\mathrm{d}}}$ and $f_{k}^{s}$ in (4b). Finally, once the distribution of friction values of dry friction, viscous friction, and every switching system is known, the anomaly classification stage produces an estimated status $\hat{\theta}$ using an SVM trained on the dataset provided. A diagram illustrating the stages of the algorithm, along with the data flow between them, is shown in Figure 5.
[…]
Figure 5: Diagram of the complete algorithm presented. The notation is explained along Section V.
IV-A Changepoint Detection
The first stage of the algorithm is the changepoint detection, which aims to determine at which steps $k$ in the signal there has been a sudden change in the total dry friction coefficient $f^{\mathrm{d}}_{k}$. This step, then, takes as input the signal from the on-board data of the satellite $\{(\omega_{k},\hat{f}_{k})\}_{k=1}^{N_{\mathrm{m}}}$ and produces as output a list of points where a sudden change of friction has been detected, defined as $\left\{k^{\mathrm{cp}}_{i}\right\}_{i=1}^{N_{\mathrm{cp}}}$, where $N_{\mathrm{cp}}$ is the number of detected changepoints.
The process of obtaining the regions of change of friction in the signal is divided into two parts. First, we compute a metric that estimates the probability that the friction has changed for every step $k$ in the data. Then, we find local maxima of this metric that are above a certain threshold. These maxima are our estimation of where the changepoints are located.
Starting with the metric, we use a normalized windowed Generalized Likelihood Ratio (wGLR) [44] with priors, defined as:
[…]
where
$\log P(X_{a:b}\ |\ f^{\mathrm{d}},f^{\mathrm{v}})$ $=\sum_{k=a}^{b}C\left[f_{k}-f^{\mathrm{d}}-f^{\mathrm{v}}\omega_{k}\right]^{2},$ (11a)
$\log P(f^{\mathrm{v}})$ $=W_{\mathrm{b}}\left(f^{\mathrm{v}}-\tilde{f}^{\mathrm{v}}\right)^{2},$ (11b)
with $C$ being a constant that gets canceled in the computation of the GLR, and $W_{\mathrm{b}}$ a parameter used to penalize large deviations of the viscous friction from the values that we expect to find in the data.
Remark 2
Like the Log-Likelihood Ratio (LLR), the metric computes the log-likelihood of two different hypotheses: a jump taking place at $k$, or no jump at all. One of the issues with the LLR is the need to consider the variables involved in both hypotheses, that is, the dry and viscous friction coefficients. This problem can be addressed through marginalization, leading to the Marginalized Likelihood Ratio or through maximization for the parameters, as we do in this problem, leading to the Generalized Likelihood Ratio [44].
The GLR was developed for cases where there is only one changepoint to detect, but there are several in our problem. Rather than computing the likelihood for the whole data, this issue is addressed by computing it in a window of $w$ points around the hypothetical changepoints. The constant $w$ is a hyperparameter of the algorithm selected from historical data, on the condition that this window size is smaller than the minimum distance between changepoints expected in the data.
The last element in the metric is the use of a prior for the viscous friction $P(f^{\mathrm{v}})$. The purpose of this element is to avoid obtaining parameters in the optimization that fit the data well but are not realistic. As we will see in Section V, a high penalization constant $W_{\textrm{b}}$ can improve the algorithm’s effectiveness when the window size $w$ is small. The base viscous friction coefficient $\bar{f}^{\mathrm{v}}$ is obtained from historical data and, as it will be explained also in Section V, does not need to be very accurate to fulfill its objective.
The purpose of this last term is to penalize parameters that deviate too much from the expected values. Its tuning is not trivial, so a parameteric study is conducted in the results section. The base viscous friction, $\tilde{f}^{\mathrm{v}}$, is defined using historical data, and should approximate the expected value of this parameter.
Once the GLR has been computed, we must determine which points correspond to changepoints. The original GLR algorithm solves this problem by taking the point $k$ at which the metric is maximized. In our case, however, we must take into account that there are several changepoints in the complete dataset. Therefore, we first find windows of contiguous measurements where the GLR is above a threshold $\text{GLR}_{\text{thr}}$ and, for each window, we designate the maximum among those points to be a changepoint $k^{\text{cp}}$. The number of changepoints will equal the number of windows above the threshold found.
The results obtained by this stage of the algorithm depend on how its main hyperparameter, the threshold, is set. Under the assumption that the dry friction coefficient is constant and the viscous friction coefficient is correctly estimated with $\tilde{f}^{\mathrm{v}}$, Wilk’s theorem shows that $\text{GLR}/\sigma_{v}^{2}\sim\chi^{2}$, where $\sigma_{v}^{2}$ is the variance of the noise in the signal [45]. This provides a good way to determine the threshold: if we want to have a false positive probability smaller than $10^{-9}$, we find the value of the threshold such that $P(\chi^{2}>\text{GLR}_{\text{thr}}/\sigma_{v}^{2})<10^{-9}$. However, a high threshold will also reduce the probability of detecting changepoints. An analysis of this trade-off is done in Section V.
The output of the algorithm is the final list $\left\{k^{\text{cp}}_{i}\right\}_{i=1}^{N_{\text{cp}}}$ of all changepoints detected in the window of measurements.
IV-B Friction Estimation
Once the changepoints have been obtained, we want to determine the dry friction coefficient between changepoints, plus the viscous friction in the data. Since the friction model we are using is linear both in the dry and the viscous friction parameters, we can use least-squares to estimate these parameters.
This stage of the algorithm takes as input the original signal $X=\left\{\left(\omega_{k},f_{k}\right)\right\}_{k=1}^{N}$, together with the location of the changepoints estimated in the previous step $\left\{k_{i}^{\text{cp}}\right\}_{i=1}^{N_{\text{cp}}}$ obtained in the changepoint detection. The algorithm’s output is a vector with the value of the dry friction coefficient at each interval, and the viscous friction, which was assumed to remain constant along the whole trajectory being analyzed.
First, we determine the regions in which the dry friction coefficient stays constant: the intervals between changepoints. Each interval $i$ is defined by the first ($k_{i}^{o}$) and last ($k_{i}^{f}$) measurement included in it. These measurement indexes are computed as:
$k_{i}^{\mathrm{o}}$ $=\begin{cases}0\quad&\text{if }i=1,\\ k_{i-1}^{\mathrm{cp}}+\Delta k_{\mathrm{error}}\quad&\text{otherwise},\end{cases}$ (12)
$k_{i}^{\mathrm{f}}$ $=\begin{cases}N\quad&\text{if }i=N_{\mathrm{cp}}+1,\\ k_{i}^{\mathrm{cp}}-\Delta k_{\mathrm{error}}\quad&\text{otherwise}.\end{cases}$ (13)
Due to the randomness in the noise affecting the measured signals, a changepoint may be detected some steps before or after the actual change in the friction coefficient. We ignore the points near the changepoints to avoid the negative effects that these may have on estimating these coefficients. We can take $\Delta k_{\mathrm{error}}=w/2$ for small window sizes, which has been observed to give good results. With this process we obtain a total of $N_{\mathrm{int}}$ intervals, where $N_{\mathrm{int}}=N_{\mathrm{cp}}+1$.
The model in (4a) can then be written for all steps $k$ in the following linear algebraic form:
$Y=H\begin{bmatrix}F\\ f^{\mathrm{v}}\end{bmatrix}+v$ (14)
where $H$ and $Y$ are defined in (15) as the concatenation of individual $H_{i}$ and $Y_{i}$ matrices (given by (16)) for each interval, $v$ is Gaussian noise with covariance $\sigma_{v}$, and $F$ is a vector containing the dry friction coefficients for each interval:
$H=\begin{bmatrix}H_{1}\\ H_{2}\\ \vdots\\ H_{N_{\text{int}}}\end{bmatrix},\quad Y=\begin{bmatrix}Y_{1}\\ Y_{2}\\ \vdots\\ Y_{N_{\text{int}}}\end{bmatrix},\quad F=\begin{bmatrix}f_{1}^{\mathrm{d}}\\ f_{2}^{\mathrm{d}}\\ \vdots\\ f_{N_{\mathrm{int}}}^{\mathrm{d}}\end{bmatrix}$, (15)
and each individual matrix is defined as:
$H_{i}$ $=\left[\begin{array}[]{ccc|c}&&&\omega_{k_{i}^{\mathrm{o}}}\\ \mathbb{0}_{\Delta k_{i}\times i}&\mathbb{1}_{\Delta k_{i}\times 1}&\mathbb{0}_{\Delta k_{i}\times(N_{\text{int}}-i)}&\vdots\\ &&&\omega_{k_{i}^{\mathrm{f}}}\\ \end{array}\right],$ (16d)
$Y_{i}$ $=\begin{bmatrix}\hat{f}_{k_{i}^{\mathrm{o}}}&\hat{f}_{k_{i}^{\mathrm{o}}+1}&\cdots&\hat{f}_{k_{i}^{\mathrm{f}}}\end{bmatrix}^{T},$ (16e)
where $\Delta k_{i}=k_{i}^{\mathrm{o}}-k_{i}^{\mathrm{f}}$ is the number of points in interval $i$, and $\mathbb{0}_{a\times b}$ and $\mathbb{1}_{a\times b}$ are matrices of shape $a\times b$ filled with zeros and ones, respectively.
Since Equation (14) is linear in $F$ and $f^{\mathrm{v}}$, we can use a least-squares to obtain an unbiased estimation of $F$ and $f^{\mathrm{v}}$:
$\begin{array}[]{rl}\begin{bmatrix}\hat{F}\\ \hat{f}^{\mathrm{v}}\end{bmatrix}&=\left(H^{T}H\right)^{-1}H^{T}Y.\end{array}$
Due to the random nature of the noise, there is an unavoidable probability that the changepoint detector is triggered at false positives. To analyze the likelihood that a certain changepoint was incorrectly detected, we define the metric $c_{i}^{\text{rjct}}$, corresponding to the increase in the squared error of the friction estimation if the changepoint $i$ is rejected (that is, ignored), and intervals $i$ and $i+1$ are assumed to have the same friction coefficient. It is computed as:
$c_{i}^{\text{rjct}}=\frac{(H^{T}H)_{i:i}(H^{T}H)_{(i+1):(i+1)}}{(H^{T}H)_{i:i}+(H^{T}H)_{(i+1):(i+1)}}(\hat{F}_{i}-\hat{F}_{i+1})^{2},$ (17)
and will be used in the next step to decide which changepoints to consider false positives.
The output of this stage is then the estimated dry friction at each interval, the estimated viscous friction, and the rejection cost of every changepoint: $(\hat{F},\hat{f}^{\mathrm{v}},\{c_{i}^{\mathrm{rjct}}\}_{i=1}^{N_{\mathrm{cp}}})$.
With the total dry and viscous friction torques at each interval obtained in this stage, the next stage will focus on determining the individual contribution of each FSS to the friction.
IV-C Maximum Likelihood Assignment
Once the total dry friction coefficient $\hat{f}^{\mathrm{d}}_{i}$ has been estimated for each interval, we are interested in decomposing this friction among the different FSSs, that is, to obtain the value of the various components of (4b) at every step, rather than only the sum of them. To achieve this, we use the knowledge of the transition probability of the subsystems defined in (5) together with the steps $k_{i}^{\mathrm{cp}}$ at which the friction changes took place to determine the most likely sequence of configurations that the subsystems went through.
Once this sequence has been estimated, we can determine the values of each subsystem’s friction coefficients $\hat{f}_{k}^{s}$ and the base dry friction $\hat{f}^{\bar{\mathrm{d}}}$. Since the dry and viscous coefficients are a function of the anomalies present in the RWA, as considered in Equation (9), we will use their values for the diagnosis.
Formally, the algorithm takes the friction values, rejection costs, and location of the changepoints as input. It estimates the base dry friction, and the friction values of each subsystem at each configuration (that is, all the elements in (4b)).
IV-C1 Maximum likelihood jumping
The first step is to determine which FSS has produced each one of the changepoints detected by the first stage of the algorithm. The output, then, is a vector $u\in\mathbb{N}^{N_{\mathrm{cp}}}$ where each entry $u_{i}$ is the index of the FSS that has caused the changepoint $i$. The objective is to determine the $u$ such that the jumps have the highest likelihood, with the probabilities given by the functions $P_{s}^{\mathrm{t}}$ and $P_{s}^{\mathrm{q}}$.
The log-likelihood of a specific sequence of jumps can be computed sequentially, which makes dynamic programming a natural choice to find the vector $u$ that maximizes it. A multistage dynamic programming is defined by four elements: a state set ($\mathcal{X}$), an input set ($\mathcal{U}$), a transition function ($h:\mathcal{X}\times\mathcal{U}\mapsto\mathcal{X}$), and a cost function ($g:\mathcal{X}\times\mathcal{U}\mapsto\mathbb{R}$) [46]. The objective is to solve the following optimization problem:
$u^{*}=$ $\operatorname*{arg\,max}_{u}\sum_{i=1}^{N_{c}}g(x_{i},u_{i})$ (18a)
s.t. $\quad x_{i+1}=h_{i}(x_{i},u_{i})$ (18b)
Extending the notation used in Section III-B, we define the state of the RWA at every interval $i$ with a tuple of two elements: the configuration $\mathbf{q}_{i}$ of each subsystem and the number of steps $\boldsymbol{\tau}_{i}$ since each subsystem changed configuration for the last time, both given as vectors with as many elements as subsystems present in the RWA, so $x_{i}=(\mathbf{q}_{i},\boldsymbol{\tau}_{i})$, with $\mathbf{q}_{i}=\begin{bmatrix}q_{i}^{1}&\cdots&q_{i}^{N_{\mathrm{s}}}\end{bmatrix}^{T}$ and $\boldsymbol{\tau}_{i}=\begin{bmatrix}\tau_{i}^{1}&\cdots&\tau_{i}^{N_{\mathrm{s}}}\end{bmatrix}^{T}$.
With this notation, $q_{i}^{s}$ represents the configuration of the FSS with index $s$ at interval $i$. Every changepoint $i$ is characterized by two variables: the number of steps since the previous changepoint ($\Delta\tau_{i}$), and the sign of the change of friction ($\Delta f_{i}$), which can be either 1-1 for a decrease or $1$ for an increase of the friction. At a changepoint, one of the subsystems changes its configuration by $\Delta f_{i}$ and its number of steps since the last changepoint is reset to 0. The other subsystems have unaffected configurations, but their $\tau$ increases by $\Delta\tau_{i}$. We choose which subsystem has jumped at each stage with the input $u_{i}$. The set of possible inputs is $\mathcal{U}\in\llbracket 1,N_{\textrm{s}}\rrbracket$, and the transition function is, according to the description given before,
$h_{i}(\mathbf{x}_{i},u_{i})$ $=\left(\mathbf{q}_{i}+\Delta f_{i}\mathbf{e}_{u_{i}},(\boldsymbol{\tau}_{i}+\Delta\tau_{i})(\mathbb{1}-\mathbf{e}_{u_{i}})\right),$ (19)
where $\mathbb{1}$ is a vector with all entries equal to $1$ and $\mathbf{e}_{u}$ is a vector with all entries 0 except the $u$-th one, which takes a value of $1$. A visualization of the possible transitions of a system is given in Figure 6.
Regarding the transition costs of the problem, they are given by the log-likelihood of the event:
$g_{i}(x_{i},u_{i})$ $=\sum_{s\not=u_{i}}\log P_{s}^{\neg\mathrm{t}}(q_{s},\tau_{s},\tau_{s}+\Delta\tau)$
$+$ $\log P_{u_{i}}^{\neg\mathrm{t}}(q_{u_{i}},\tau_{u_{i}},\tau_{u_{i}}+\Delta\tau_{u_{i}}-1)$ (20)
$+$ $\log P_{u_{i}}^{\mathrm{t}}(q_{u_{i}},\tau_{u_{i}}+\Delta\tau)+\log P_{u_{i}}^{\mathrm{q}}(q_{u_{i}}+\Delta f_{i}|q_{u_{i}}).$
In this computation, we define $P_{s}^{\neg t}(q,k_{1},k_{2})$ as the probability of the switching system $s$ not transitioning away from its configuration $q$ between steps $k_{1}$ and $k_{2}$. It can be calculated as
$P_{s}^{\neg\mathrm{t}}(q,k_{1},k_{2})=\sum_{k=k_{1}}^{k_{2}}\log\left[1-P_{s}^{\mathrm{t}}(q,k)\right].$ (21)
Recalling Remark 1, we know $P_{s}^{\mathrm{t}}(q,\tau)$ and $P_{s}^{\mathrm{q}}(q^{\prime}|q)$, so we can compute all the log-likelihood scores, either as the dynamic programming problem is being solved, or in advance, and then use look-up tables during the execution of the algorithm. The objective, then, is to obtain a sequence of inputs $u$ maximizes log-likelihood. This is a standard dynamic programming problem that can be solved with classical algorithms [46]. Note that while the possible values of $\mathbf{q}$ are relatively limited, this is not the case for $\boldsymbol{\tau}$, and the number of possible states may increase exponentially with the number of stages.
For the initial state $x_{0}$ of the dynamic programming problem, we take an optimistic approach: for $\mathbf{q}_{0}$, we initialize the algorithm considering all possible values, and keep the one that provides the maximum final score. Regarding $\boldsymbol{\tau}_{0}$, we select the value that maximizes the score in (20).
The solution $u^{*}$, obtained by solving (18), provides then the index of FSS that produced each sudden change of friction observed in the data. Additionally, we denote $\mathbf{x}^{*}$ the sequence of states obtained by applying this input.
$\mathbf{q}_{i}$ $=\begin{bmatrix}0\\ 1\end{bmatrix}$ $\boldsymbol{\tau}_{i}$ $=\begin{bmatrix}50\\ 0\end{bmatrix}$ $\mathbf{q}_{i+1}=\begin{bmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}\\ 1\end{bmatrix},\boldsymbol{\tau}_{i+1}=\begin{bmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}\\ 80\end{bmatrix}$ $\mathbf{q}_{i+1}=\begin{bmatrix}0\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2}\end{bmatrix},\boldsymbol{\tau}_{i+1}=\begin{bmatrix}130\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}\end{bmatrix}$ $\mathbf{q}_{i+1}=\begin{bmatrix}0\\ 1\end{bmatrix},\boldsymbol{\tau}_{i+1}=\begin{bmatrix}130\\ 80\end{bmatrix}$ $u_{i}=1$$u_{i}=2$$u_{i}=0$
Figure 6: Example of possible transitions during the dynamic programming problem, given that the next changepoint has $\Delta f_{i}=+1$, and $\Delta\tau_{i}=80$.
IV-C2 Changepoint rejection
The algorithm presented is inherently fragile: if a changepoint is missed during the first stage of the algorithm, or a false positive is obtained, the algorithm will fail to obtain the correct sequence of configurations of the FSSs. Unfortunately, designing a changepoint detection algorithm that detects all actual changepoints with no possibility of false positives is impossible due to the randomness of the noise that affects the friction torque measurements.
Since we want the algorithm to perform well even in the presence of false positives, we robustify it by including in the dynamic programming problem the option at every stage of ignoring the changepoint detected, rather than fitting it in one of the FSS. This is done by including 0 as an additional possible input, shown in Figure 6, with the following transition and cost functions:
$h_{i}(x_{i},0)$ $=\left(q_{i},\tau_{i}+\mathbb{1}\Delta\tau_{i}\right).$ (22)
The cost of this transition is given by the log-likelihood of none of the switching systems jumping during this stage, plus the added cost of rejecting this changepoint. Formally, this is written as
$g(x_{i},0)$ $=c_{i}^{\text{rjct}}+\sum_{s}P^{\neg\mathrm{t}}(q_{s},\tau_{s},\tau_{s}+\Delta\tau).$ (23)
Recalling (17), this additional cost is given by how much the squared error of the regression increases when the changepoint is removed and the same friction coefficient is estimated before and after it. Therefore, the algorithm is more likely to ignore changepoints when the intervals before and after them have similar friction coefficients.
IV-C3 Friction Values
Once the dynamic programming algorithm is finished, we get a vector $u$ containing the index of the FSS that changed configuration at each changepoint and an array $\mathbf{q}$ with the configuration of every FSS after every changepoint. The next objective is to estimate the base dry friction $\hat{f}^{\bar{\mathrm{d}}}$ and, for each FSS, the sequence of configurations and friction values it goes through. We denote them as $\mathcal{Q}_{s}$ and $\mathcal{F}_{s}$ respectively, both defined as lists of values in the programming sense, where $s$ denotes the index of the FSS. At every interval, the total dry friction torque must be the base dry friction plus the contribution of each FSS as given by (4b).
This problem is solved in Algorithm 1. We start by assuming that, before the first changepoints, the friction produced by each FSS is equal to 0. Then, we compute the following friction values from the differences obtained at each changepoint. Finally, since the minimum value produced by an FSS is assumed to be 0, we adjust the coefficients to get the final values.
[…]
7:for $i\leftarrow 2$, $N_{\text{int}}$ do
8:  $\triangleright$ At every changepoints, add the $f$ and $q$ to the corresponding FSS.
9:  $\mathcal{F}_{u_{i}}$.append($\mathcal{F}_{u_{i}}[-1]+\hat{f}_{i}-\hat{f}_{i-1}$)
10:  $\mathcal{Q}_{u_{i}}$.append($q_{i}$)
11:end for
12:for $s\leftarrow 1$, $N_{s}$ do
13:  $\triangleright$ Adjust $\mathcal{F}_{s}$ so its minimum is equal to 0
14:  $f_{\text{min}}=\min\mathcal{F}[s]$
15:  for $i\leftarrow 1$, $\text{len}(\mathcal{F}[s])$ do
16:   $\mathcal{F}_{s}[i]\leftarrow\mathcal{F}_{s}[i]-f_{\text{min}}$
17:  end for
18:  $f_{\text{base}}\leftarrow f_{s}+f_{\text{min}}$
19:end for
Algorithm 1 Obtain $\mathcal{Q}_{s}$ and $\mathcal{F}_{s}$
The final output of this stage is a tuple of four elements: the estimated base dry friction ($\hat{f}^{\bar{d}}$, the estimated viscous friction coefficient ($\hat{f}^{v}$), and the sequence of states ($\mathcal{Q}_{s}$) and friction values ($\mathcal{F}_{s}$) each FSS goes through: $(\hat{f}^{\bar{\mathrm{d}}},\hat{f}^{\mathrm{v}},\{\mathcal{Q}_{s},\mathcal{F}_{s}\}_{s=1}^{N_{\mathrm{s}}})$.
IV-D Anomaly Classification
The final stage is the anomaly classification, in which the friction coefficients obtained in the last step are used to determine the anomaly status of each component affecting the friction torque in the RWA.
This step, then, takes the assigned friction coefficients $\hat{f}^{\bar{d}}$, $\hat{f}^{v}$ and $\left\{(\mathcal{Q}_{s},\mathcal{F}_{s})\right\}$ for each FSS $s$, and produces a diagnosis for the dry friction $\hat{\theta}_{d}$, a diagnosis for the viscous friction $\hat{\theta}_{v}$ and one diagnosis for each switching system $\hat{\theta}_{s}$. These diagnoses are independent from each other: $\hat{\theta}_{d}$ depends only on $\hat{f}^{\bar{d}}$, $\hat{\theta}_{v}$ on $\hat{f}^{v}$, and the diagnosis of each subsystem is done with the friction values associated to it. We can then treat each of these classification problems as independent.
As mentioned in Section III, we have knowledge of which part of the friction model is affected by each anomaly; however, we do not know how it is affected. To address this shortcoming, we use the labeled dataset $D$, where we have a collection of datapoints with the anomalies associated with them. For each datapoint in the dataset, we run the first three steps of the algorithm to obtain a processed dataset
$D_{\mathrm{p}}$ $=\left\{\left(\hat{f}^{\bar{\mathrm{d}}},\hat{f}^{\mathrm{v}},\left\{\left(\mathcal{Q}_{s},\mathcal{F}_{s}\right)\right\}_{s=1}^{N_{\mathrm{s}}}\right)\right\}_{i=1}^{N_{\mathrm{D}}}.$ (24)
This dataset we can be split into $2+N_{s}$ different datasets, one for each anomaly:
$D^{\mathrm{\bar{d}}}$ $=\left\{\left(\hat{f}^{\bar{\mathrm{d}}},\theta^{\mathrm{d}}\right)\right\},\qquad D^{\mathrm{v}}=\left\{\left(\hat{f}^{\mathrm{v}},\theta^{\mathrm{v}}\right)\right\},$ (25a)
$D^{s}$ $=\left\{\left(\left\{\left(\mathcal{Q}_{s},\mathcal{F}_{s}\right)\right\}_{s=1}^{N_{\mathrm{s}}},\theta^{s}\right)\right\}\forall s\in[1,N_{s}].$ (25b)
In our problem, the status of each anomaly is given by a simple binary label: active or inactive. In our pursue of a simple and effective tool for the classification, we choose to use Support Vector Machines (SVM) [47]. By defining one SVM per anomaly, we can train them on the corresponding independent dataset.
An SVM is a classification method that takes as input a vector $z\in\mathbb{R}^{n}$ and produces an output $y\in\{-1,1\}$ according to the following equation:
$\hat{\theta}(h)=\text{sign}\left(w^{T}z+b\right).$ (26)
The classifier is parametrized by the weights $w$ and the bias $b$. Although not as versatile as state-of-the-art neural networks, SVMs have two important advantages due to their reduced number of parameters: they are less prone to overfitting, and doing inference is computationally less expensive. Although we use SVMs to classify all anomalies, the implementations vary depending on whether we are determining a dry friction anomaly, a viscous friction anomaly, or an anomaly related to an FSS.
IV-D1 Dry and viscous friction anomalies
The first two anomalies involve an increased value in the base dry and viscous friction coefficients, respectively. The classification, therefore, consists in analyzing a single value in each case, either $z=\hat{f}^{\bar{\mathrm{d}}}$ or $z=\hat{f}^{\mathrm{v}}$.
This is accomplished by defining an SVM that takes values in $\mathbb{R}^{1}$ as inputs, making the classification a simple threshold learned from the dataset.
IV-D2 Subsystem anomalies
Classifying anomalies for switching systems is more challenging, since we do not have a single value to classify with, but a sequence of them in the form of $\mathcal{F}_{s}$. As mentioned in Subsection III-C, an anomaly impacts the probability distribution of the friction produced by the FSS, so constructing a histogram from the friction values and feeding it to an SVM that does the classification is a reasonable approach.
The normalized histogram vector of the FSS with index $s$ is defined as $z\in\mathbb{R}^{n_{\text{bins}}}$, obtained from a sequence of friction values $f_{k}$ as
$\forall j\in\{1,\dots,n_{\text{bins}}\},\quad z_{j}=\frac{1}{|\mathcal{F}_{s}|}\sum_{f\in\mathcal{F}_{s}}\mathbb{1}_{f\in[r_{j},r_{j+1}]},$ (27a)
$r_{j}=r_{\text{min}}+\frac{j}{n_{\text{bins}}}(r_{\text{max}}-r_{\text{min}}).$ (27b)
Two hyperparameters must be defined, the number of bins ($n_{\text{bins}}$) and the limit values ($r_{\text{min}}$ and $r_{\text{max}}$). The latter can be determined from the historical data, by annotating the minimum and maximum values of friction observed. The number of bins to select is a trade-off between accuracy and complexity. A small number of bins may make distinguishing an anomaly impossible, while a large number will lead to a large histogram vector, which may complicate training and lead to overfitting. For our task, we select a value of $20$ bins, and later analyze the effect of this hyperparameter on the accuracy of the algorithm in Section V-D2.
The last decision to take is which values of friction to include in the histogram. From the previous section, we recall that we obtain, for each subsystem, a collection of friction values paired with their corresponding configurations. Based on our knowledge of the anomaly, we can decide to include all friction values produced by the FSS in the histogram, or only those related to specific configurations (for example, only friction values obtained with $q=1$).
The output of this final stage is a tuple with the status of each possible anomaly, $(\hat{\theta}^{\mathrm{d}},\hat{\theta}^{\mathrm{v}},\{\hat{\theta}^{s}\}_{s=1}^{N_{\mathrm{s}}})$.
V Results
We now proceed to analyze the model and the algorithm developed in the paper. As mentioned in Section I, this work aims to create an anomaly detection algorithm for the TAS use case.
There are three objectives pursued in this section. The first is to show the adequacy of the model proposed in Section III to the dataset provided by TAS, so we can confirm that it captures the behavior observed in the dataset with accuracy. The second objective is to analyze the algorithm’s effectiveness in detecting anomalies for the TAS use case. Finally, we want to study the effect of some hyperparameters of the algorithm whose tuning is not straightforward.
The algorithm’s structure, composed of four sequential stages, is beneficial for analysis, as each part of the algorithm can be analyzed independently to ensure it works as expected. This way, our analysis is not limited to the final accuracy obtained by the last stage. Unfortunately, while we have a ground truth in the classification because the dataset is labeled, we do not have such a reference for the output of the different stages of the algorithm. Therefore, we will use various methods to assess other aspects of the work.
Rather than organizing this section by objectives, we organize it by stages of the algorithm. We start by analyzing the changepoint detection stage using synthetic data. For the friction estimation and changepoint assignment, we use the TAS dataset to analyze some performance metrics so, even without a ground truth to compare to, we can study the behavior of these stages. Finally, we will use the labels provided in the TAS dataset to assess the accuracy of the complete algorithm at determining the anomaly present in the RWA. Additionally, we investigate the computational cost of the whole process to establish its suitability for running in a low-power device located in the satellite.
V-A Changepoint detection
We analyze the first part of the algorithm, which takes a time-series $\{(\omega_{k},f_{k})\}_{k=1}^{N_{\mathrm{m}}}$ and determines the position of the multiple abrupt changes of friction.
In the analysis of a changepoint detection algorithm, there are two main metrics of performance: the Missed Detection Rate (MDR), defined as the probability of not detecting a changepoint, and the Average Run Length (ARL), defined as the average number of points the algorithm goes through before detecting a false changepoint in the data [44]. Since it is always possible to improve one at the expense of the other (by decreasing or increasing the threshold $\text{GLR}_{\text{thr}}$), the challenge lies in achieving a high ARL with a low MDL.
To assess the performance of the changepoint detection, it is important to remember how the subsequent stages of the algorithm process changepoints. As mentioned in Section IV-C2, an undetected changepoint can invalidate the algorithm’s output while a false positive can be rejected during the assignment stage. This implies that while high ARL is beneficial, as it will lead to fewer false positives, a low MDR is fundamental for the algorithm’s success.
For our analysis, we will fix the MDR to $0.1\%$ and determine the threshold value at which such sensitivity is achieved using the $\chi^{2}$ distribution. Then, we obtain the average run length with this threshold, which we call the $\text{ARL}^{0.1\%}$. A satisfactory value of ARL depends on the tolerance of the algorithm to false positives, and the number of points we process in a single window of measurements. In the TAS case, we analyze windows of around $80000$ measurements, so an ARL of around $10^{6}$ is a good target if we do not want to get more than $1$ false changepoint per window. Since the TAS dataset has no annotated changepoint locations, these results are obtained using our own simulations, in which we can place changepoints in a controlled way.
V-A1 Bias selection
First, we analyze the effect of bias on the algorithm’s performance. Recalling Subsection IV-A, $W_{\mathrm{b}}$ is a tunable parameter, representing the weight of the prior on the viscous friction coefficient in (10). Since there is no clear way of choosing this value to obtain better results, we run a parametric study to observe how it affects the $\text{ARL}^{0.1\%}$ for different window sizes, assuming that the changes of dry friction we are trying to detect have a magnitude of $3$ times the noise standard deviation. The results are shown in Figure 7, left plot. A higher window size $w$ leads to better algorithm results, as the effect of the noise is reduced when we use more points the w-GLR. However, increasing $w$ is not always feasible, because it is limited by the minimum distance we expect to see between changepoints. We can observe how the $W_{\mathrm{b}}$ improves the algorithm’s performance for different window sizes. As an example, using a window size of $20$, we can improve the $ARL^{0.01\%}$ from $10^{4}$ to $10^{10}$ by increasing $W_{\mathrm{b}}$ from $10^{-8}$ to $10^{-4}$. The bias term is, then, a good way of improving the performance of the changepoint detection when the window size is constrained, and it seems from this first analysis that $W_{\mathrm{b}}$ should be set to be arbitrarily large.
However, in the previous plot, we assumed that our estimated friction $\tilde{f}^{\mathrm{v}}$ is equal to the actual viscous friction $f^{\mathrm{v}}$. Unfortunately, this may not be the case, since the latter is not known in advance. Considering the effect of the bias $\tilde{f}^{\mathrm{v}}-f^{\mathrm{v}}$ is more complex: when the error is not zero, the w-GLR no longer follows a $\chi^{2}$ distribution. We can, however, estimate the effect of this error to be in the order of $W_{\mathrm{b}}(\tilde{f}^{\mathrm{v}}-f^{\mathrm{v}})^{2}$. As the original distribution of the w-GLR is of order unit, we should ensure that $W_{\mathrm{b}}(\tilde{f}^{\mathrm{v}}-f^{\mathrm{v}})^{2}\lll 1$ so the error does not have a noticeable effect. Using $W_{\mathrm{b}}=10^{-4}$, this implies that we need $|\tilde{f}^{\mathrm{v}}-f^{\mathrm{v}}|\lll 10^{2}$, an easily achievable condition given that the viscous friction coefficient is of order $1$.
This gives a simple rule for tuning $W_{b}$: if the window size is large enough (more than 50 points), it can be set to 0. If not, a high value can be set to improve the results, but keeping in mind that if the prior of the viscous friction is not reliable, the performance may degrade and produce too many false positives.
V-A2 Changepoint properties
We now analyze the effect of the window size $w$ and the size of the changepoints $\Delta f_{\mathrm{cp}}$ on the $\text{ARL}^{0.1\%}$, while keeping $W_{b}=10^{-4}$. The results are shown in Figure 7, right plot. We can see that better results are achieved with either high window sizes or changepoint strengths. This shows a basic limitation of the algorithm: if we want to detect small friction changes, we need to use large window sizes. However, this might not be possible if the changepoints that we want to detect may be too close together, in which case we may need to increase the value of $W_{\mathrm{b}}$.
Figure 7: Left: value of $\text{ARL}^{0.1\%}$ for different window sizes $w$ and bias $W_{\mathrm{b}}$, assuming changes in friction of $\Delta f_{\text{cp}}=3\sigma_{v}$. Right: effect of window size and friction change on $\text{ARL}^{0.1\%}$, with $W_{b}=10^{-4}$.
As a last note, the average run lengths obtained in these analyses assume that the data contains no abnormalities, that is, the friction is just a combination of dry and viscous friction and Gaussian noise as given by (4a). Real data may contain imperfections that trigger additional, unavoidable false changepoints. Therefore, a high ARL is necessary for reducing the number of false changepoints, but it does not guarantee it.
V-B Friction Estimation
We now analyze the output of the second step of the algorithm: the friction estimation.
This algorithm step is of no significant interest since the Least-Squares algorithm has well-known properties, and the only parameter to adjust is $\Delta k$, which does not have a large impact on the output. We can, on the other hand, analyze how well the friction coefficients $\hat{f}^{\mathrm{d}}_{i}$ and $\hat{f}^{v}$ obtained in this stage approximate the friction $\hat{f}_{k}$ in the data.
A small error is an indicator of a good fit of the model to the data (at least when it comes to (4a), since we are not considering the different FSM yet) and a correct estimation of the changepoint locations. This is because, if changes in the friction were not detected, the algorithm would try to fit the same coefficient to intervals of measurements with different dry friction coefficients, which would produce a high error.
We use the Root Mean Square Error (RMSE) as a metric of fit, which measures the error between the friction values in the data and those estimated using the intervals and the linear regression run in stages I and II of the algorithm. The RMSE for a given time series is computed as
RMSE $=\sqrt{\frac{1}{N_{\mathrm{m}}}\sum_{i=1}^{N_{\mathrm{int}}}\left[\sum_{k=k_{i}^{o}}^{k_{i}^{f}}(\hat{f}_{k}-\hat{f}^{\mathrm{d}}_{i}-\omega\hat{f}^{\mathrm{v}})^{2}\right]}.$ (28)
Additionally, we compute for each time-series in the TAS dataset a “naive” RMSE, obtained by ignoring changepoints and assuming the dry friction coefficient to be constant along the whole time-series. The histograms of both metrics are shown in Figure 8, together with a 2D histogram that relates both RMSEs. To take into account the effect of the noise on the error, we estimate its standard deviation using a short window of data and subtract it from the RMSE, resulting in what we call the excess RMSE.
Figure 8: Excess RMSE in the datapoints of the TAS dataset. The top histogram is obtained by computing a single dry and viscous friction coefficient for the complete time-series, the right one by obtaining several dry friction coefficients from steps 1 and 2 of the algorithm. The mean values of the histograms are $0.13$ and $0.0025$, respectively.
While the excess RMSE obtained under the assumption of constant $f^{d}$ has an average value of $0.13$, it goes down to $0.0024$ when considering a variable dry friction coefficient. This considerable decrease reveals the importance of considering the jumps of the dry friction coefficient. The 2D histograms, on the other hand, show a lack of correlation between both errors. This suggests that no other significant phenomena were present in the data that were not accounted for, as it would have produced a similar increase in RMSE with single and variable dry friction coefficients.
Although the RMSE obtained after accounting for changepoints is small enough, considering that the minimum friction observed in the data is around $1.0$, it is higher than what we would expect if the noise were the only source of error. If such were the case, we would expect a mean excess RMSE of 0 with a very small variance.
To investigate further, we collect the absolute error obtained for every measurement in the TAS dataset and generate a survival function, shown in Figure 9. Assuming that Gaussian noise is the only source of error, we can compute the theoretical survival function using the $\chi^{2}$ distribution and compare it to the experimental one. Differences between the two distributions give information about errors due to incorrect estimations of the friction coefficients.
Figure 9: Survival function of the absolute error of the friction estimation for all measurements of all datapoints in the dataset, together with the theoretical distribution if random noise was the only cause of error.
We observe a very close match between both functions until an error of approximately $0.08$. As the value of the experimental survival function is approximately $10^{-3}$, we can see that out of every $1000$ points, $999$ have an error that falls in the expected distribution, suggesting a good fit of the data. However, we observe many points with a higher error than expected, explaining the high mean RMSE observed in Figure 8. The fact that the fraction of points with high error is so small suggests irregularities in the data and occasional estimation errors, rather than some fundamental issue in the model or the algorithm.
As in the previous subsection, there is a word of caution. Although a low RMSE is a good indicator of the correct fit of the coefficients, it is not a flawless metric. An algorithm that detects more changepoints than necessary will produce a reduced RMSE. Although the previous subsection showed that the false positive detection was not a cause of concern, the following subsection will explore the issue of false positives in more detail.
V-C Assignment
We analyze the assignment stage of the algorithm, tasked with determining which changepoints are caused by which FSS by running a dynamic programming algorithm to maximize the likelihood of the assignment. As with the case of the changepoints, we do not have a ground truth of what the correct assignment of changepoints is. Instead, we will focus on two elements of the algorithm: the number of changepoints rejected, and the number of iterations required before an assignment is obtained.
Table II shows the number of changepoints rejected as a histogram taken over all the datapoints in the TAS dataset. An unusually low number of rejections can be observed: 62% of the datapoints are processed with no changepoints being rejected, while 96% of the runs require three or fewer rejections. This shows that, while rejection is a needed component of the algorithm, it does not seem to be abused, considering that the average number of changepoints per run is around 26 (with a standard deviation of 10).
N rejections $\mathbf{1}$ $\mathbf{2}$ $\mathbf{3}$ $\mathbf{4}$ $\mathbf{5}$ $\mathbf{6}$ $\boldsymbol{\geq}\mathbf{7}$
TAS runs $\%$] $62.6$ $18.3$ $12.2$ $4.5$ $2.4$ $0.5$ $0.7$
TABLE II: Percentage of runs in the TAS dataset by number of changepoints rejected.
We are also interested in the computational cost of the algorithm, since it is well known that dynamic programming does not scale well with the number of states [46]. So, it is essential to remember this when analyzing the algorithm. To do this, we focus on the number of iterations the algorithm takes to reach an optimal assignment, as it is proportional to the time it takes the algorithm to finish. In Figure 10, we show a scatter plot of runs from the TAS dataset by the number of changepoints they had, and the number of iterations it takes to compute their assignment. Additionally, we distinguish three groups according to the number of rejected changepoints: green for no rejections (accounting for $62\%$ of the points), blue for 1 to 3 rejections ($34\%$ of the points), and red for more than three rejections ($4\%$ of the points).
Figure 10: Relation between the number of iterations run by the assignment algorithm before obtaining the maximum likelihood assignment, and the number of changepoints to assign. The mean of $116.6$ iterations was obtained across all datapoints in the TAS dataset.
For the case of no rejections, we observe the number of iterations to scale roughly linearly with the number of changepoints. However, some outliers can be observed (likely coming from time-series in which several assignment hypotheses were kept until the end). However, this linear scaling is not as straightforward when the number of rejected changepoints increases. In this situation, the relation between the number of changepoints and the number of iterations becomes less predictable: at best, it stays linear, while at worst, it gets exponential. This randomness is suspected to be due to where the rejected changepoint is located. A false changepoint may be easy to detect, or require ruling out many alternative hypotheses, notably increasing the number of iterations.
This points out the main limitation of rejection of changepoints, as it makes the assignment problem much more expensive. We should, therefore, be careful with the Average Run Length set for the changepoint detection, and not expect the assignment to take care of all false positives.
As a final note, we must remember that, on average, the number of iterations is relatively small, around 115 iterations. This is because, although this number can reach high values when the number of changepoints and/or rejections is high, both numbers are relatively small for most runs.
V-D Classification
We present the paper’s main results now, namely, the algorithm’s accuracy in determining the anomaly affecting the RWA given a time series of measurements of spin rate and friction torque captured by the onboard sensors. Since the TAS dataset had a ground truth of the anomaly affecting the simulated RWAs used to generate the datasets, we can easily assess how well the algorithm performs the detection and classification of the anomaly.
Note that this step represents the combined work of every previous algorithm step. Hence, a good accuracy is itself an indicator of the capability of the complete algorithm to analyze the onboard data and determine the anomaly or anomalies affecting the RWA. In addition to the accuracy of the diagnosis, we study the impact of the number of bins used in the histogram generation on anomalies related to the FSS.
V-D1 Anomaly classification
To determine how accurate the algorithm is at determining the anomaly present in an RWA, we compare its output to the labels in the TAS dataset. Although the algorithm can determine several anomalies at once, assigning a binary label to each one, time-series in the dataset can only have one anomaly at most. We consider the five possible actual statuses (1 nominal + 4 anomalous), according to the label, and determine in what percentage of the data points with that status the algorithm detected each anomaly.
The classification step involves training an SVM, done by splitting the dataset into a training set (20% of the datapoints) and a validation set (80% of the datapoints). Since this introduces some randomness in the process, we run the training and validation 30 times and show the minimum, mean, and maximum values to verify the stability of the classification. The results are shown in Figure 11.
Figure 11: Anomaly detection probabilities obtained with 30 random splits of training and validation sets. The top number is the highest, the middle is the mean, and the bottom is the lowest among the 30 tests. Note that percentages do not add up to 100%, since the same time-series can be classified with multiple anomalies simultaneously.
For the four anomalies considered, the probability of them being correctly detected is around $95\%$. The only exception is the FSS 2, which has an accuracy of $100\%$. Perfect accuracy in the classification implies that a separating plane exists between the histograms of the friction values of the FSS 2 with and without the anomaly.
Another point of consideration is the probability that a datapoint is misclassified. These probabilities are generally low, with the highest number found in the probability that an anomaly in FSS 2 is detected for a datapoint with an anomaly in FSS 1, at $14\%$. This can be attributed to some confusion in the assignment process, which assigns a changepoint produced by one FSS to a different one.
Finally, we observe that the differences between the lowest and highest values obtained in different trainings are very small, with the most significant variation around $5\%$. This can be attributed to the simplicity of the SVM, since a small number of tunable hyperparameters reduces the risk of overfitting.
V-D2 Number of bins
For the TAS dataset, we used 40 bins to generate the histograms used to classify anomalies of both switching systems. To support this decision, we tested different bin numbers ($n_{\mathrm{bins}}$) and computed the resulting train and validation loss in Figure 12.
Figure 12: Train and validation loss with varying number of bins for classification of anomalies of subsystem 1.
The plot shows an expected decrease in the train loss as the number of bins increases, since the dimension of the classification increases with it. However, many bins lead to overfitting, as training captures only the specific shape of the datapoints used in training, and the validation loss increases. This shows a good level of robustness of the algorithm to the number of bins, which can again be attributed to the use of SVMs for the classification.
V-E Computation Time
We now evaluate the times it takes the algorithm to run on the datapoints of the dataset provided by TAS, each of them containing around 80000 measurements of spin rate and friction. The tests were run on a laptop with an Intel® Core™ Ultra 7 165U.
We first analyze the time taken by each step of the algorithm to run, together with the total computation time. This results in Figure 13, a violin plot with logarithmic scale to represent the large differences in computation time of the different steps.
Figure 13: Distribution of computation time taken by each step of the algorithm, and by the complete algorithm. Each violin plot has 3 horizontal lines corresponding to the maximum, the mean, and the minimum times.
The classification step takes less than 10 ms to run, which makes its contribution negligible. This is to be expected: generating the histograms is relatively cheap, and running an SVM involves little more than an inner product, which makes this step the cheapest to run, with a mean and maximum times of 1 and 10 ms respectively (the variations is due to a higher number of changepoints leading to more values to be included in the histograms). The changepoint detection is the most expensive step to run, with an average runtime of 100 ms, but its cost has little variation as the computation of the LLR is done using vector operations.
The assignment component exhibits the most interesting behavior: it is fairly cheap on average (around 30 ms to run), but can reach up to 3 seconds in the dataset tested. The observations done on Figure 10 suggest that such increases are related to a high number of changepoints rejected, which were observed to cause an exponential increase in the number of iterations of the algorithm. Although such an increase is very negative for the performance of the whole algorithm, only a very small percentage (around 5%) of the datapoints had more than 3 rejected changepoints. In any cases, the computational times are very reasonable given the large number of measurements that each time-series contains, which makes the algorithm suitable to be deployed in the satellite.
VI Conclusion
We presented an algorithm to determine the anomaly present in an RWA from a window of measurements of spin rate and friction torque. The algorithm has 4 steps in order to address the peculiarities of the data, dealing with the need to detect changepoints and the properties of each anomaly shown exclusively through a labeled dataset. The results show an overall satisfactory performance of the algorithm, with an accuracy in the detection of anomalies of around $95\%$. Although the algorithm was developod with the TAS use case in mind, it can be applied to other friction systems that show a jumping behavior in their dry friction coefficient.
The analysis of the different steps of the algorithm provides an insight into possible future research on the topic. For instance, the changepoint algorithm could be improved using variable window sizes instead of fixed, although this would come at an additional computational cost. The assignment algorithm, although functional, can be improved using an improved algorithm developed for dynamic programming. Another aspect to consider is the need to know the transition functions of the FSS ($P_{s}^{\mathrm{t}}$) in order to run the algorithm: automatic methods for identifying them would make the algorithm easier to apply to new systems, without the need to manually identify them.
Appendix A Biography Section